# Learning Robust Anymodal Segmentor with Unimodal and Cross-modal Distillation

### v1

@@ F E L FE FL EL FEL
CMX (Zhang et al., 2023a) T-ITS 2023 2.52 2.35 3.01 41.15 41.25 2.56 42.27 19.30
CMNeXt (Zhang et al., 2023b) CVPR 2023 FEL 3.50 2.77 2.64 6.63 10.28 3.14 46.66 10.80
MAGIC (Zheng et al., 2024b) ECCV 2024 43.22 2.68 22.95 43.51 49.05 22.98 49.02 33.34
Any2Seg (Zheng et al., 2024a) ECCV 2024 44.40 3.17 22.33 44.51 49.96 22.63 50.00 33.86
@@
Cross-modal Correspondence Distillation.
To address this, we leverage cross-modal correspondences between “easy-to-learn” and “hard-to-learn” modalities to achieve a more balanced performance. By distilling these cross-modal relationships from the teacher model to the student model—referred to as the “anymodal” segmentor—we effectively mitigate unimodal bias and enhance performance across all modalities.
The distillation of cross-modal knowledge between student features $\{g_{r}^{i},g_{d}^{i}\}_{i=1}^{4}$ and teacher features $\{f_{r}^{i},f_{d}^{i},f_{e}^{i},f_{l}^{i}\}_{i=1}^{4}$ is achieved through:
$\mathcal{L}_{cmd}=\sum_{i=1}^{4}\sum_{j=1}^{C_{i}}\mathcal{S}\left(g_{d}^{i,j},g_{r}^{i,j}\right)\log\left(\frac{\mathcal{S}\left(g_{d}^{i,j},g_{r}^{i,j}\right)}{\mathcal{S}\left(f_{d}^{i,j},f_{r}^{i,j}\right)}\right),$ (5)
where $\mathcal{S}(x,y)=\frac{x\cdot y}{||x||||y||}$ denotes cosine similarity between feature vectors $x$ and $y$. This formulation aligns cross-modal representations by minimizing the discrepancy between teacher and student model similarities across modalities. After applying cross-modal correspondence distillation, the inherent unimodal bias in this task—as well as the bias introduced by unimodal knowledge distillation—are largely mitigated, as shown in Tab. 5.
@@ 3.3 Modality-agnostic Distillation
The total loss for training the anymodal student segmentor combines the modality-agnostic distillation loss and the supervised loss as follows:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{sup}+\lambda\mathcal{L}_{mad}+\alpha\mathcal{L}_{umd}+\beta\mathcal{L}_{cmd},$ (8)
where $\lambda_{\text{mad}}$ and $\lambda_{\text{sup}}$ are weighting factors that balance the contributions of the two loss terms.
Table 2: Results of anymodal semantic segmentation validation with three modalities on synthetic benchmark DELIVER dataset using SegFormer-B0 as backbone model.
@@
CMNeXt (Zhang et al., 2023b) 0.86 0.49 0.66 0.37 47.06 9.97 13.75 2.63 1.73 2.85 59.03 59.18 14.73 59.18 39.07 20.77
MAGIC (Zheng et al., 2024b) 32.60 55.06 0.52 0.39 63.32 33.02 33.12 55.16 55.17 0.26 63.37 63.36 33.32 55.26 63.40 40.49
@@
Table 3: Ablation study of different loss function combinations in our framework on MUSES dataset (Brödermann et al., 2024).
@@
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our proposed method on both synthetic and real-world multi-sensor datasets. The MUSES dataset (Brödermann et al., 2024) consists of driving sequences recorded in Switzerland, specifically designed to tackle challenges posed by adverse visual conditions. It features multi-sensor readings, including a high-resolution frame camera, an event camera, and MEMS LiDAR, providing complementary modalities to enhance annotation quality and support robust multimodal semantic segmentation. Each sequence is meticulously annotated with high-quality 2D panoptic labels, ensuring accurate ground truth for comprehensive benchmarking. The DELIVER dataset (Zhang et al., 2023b) is a diverse multimodal dataset comprising RGB, depth, LiDAR, and event data across 25 semantic categories. It covers various environmental conditions and includes scenarios with sensor failures, enabling thorough evaluations under challenging situations. For our experiments, we follow the official data processing and split protocols.
Implementation Details. All experiments on the MUSES dataset were conducted on 8 NVIDIA 3090 GPUs, while experiments on the DELIVER dataset utilized 4 NVIDIA A100 GPUs. The initial learning rate was set to $6\times 10^{-5}$ and adjusted using a polynomial decay strategy with a power of 0.9 over 200 epochs. Additionally, a 10-epoch warm-up phase was applied at 10% of the initial learning rate to stabilize training. The AdamW optimizer was employed, and the effective batch size for both datasets was set to 16. Input modality data was cropped to $1024\times 1024$ resolution for consistency across benchmarks.
4.2 Results
As presented in Tab. 1, the proposed method achieves the highest mIoU of 40.23, outperforming all state-of-the-art (SoTA) methods. CMX (Zhang et al., 2023a) and CMNeXt (Zhang et al., 2023b) show weaker performance (mIoU: 19.30 and 10.80, respectively), struggling with individual modalities due to their reliance on RGB. Similarly, MAGIC (Zheng et al., 2024b) and Any2Seg (Zheng et al., 2024a) rely heavily on depth data, resulting in significant performance drops when this modality is unavailable, revealing their unimodal bias. In contrast, our method demonstrates superior performance across all metrics, effectively mitigating unimodal bias. It achieves strong results on individual modalities—RGB (46.01), Event (19.57), and LiDAR (32.13)—and excels in paired modalities, such as FE (46.29), FL (51.25), and EL (35.21), as well as combined modalities like FEL (51.14), showcasing its cross-modal learning capability. Notably, our method achieves significant gains in challenging scenarios, such as Event data (+16.40) and paired modalities involving LiDAR (e.g., FL: +12.58), highlighting its ability to leverage multimodal data and handle complex interactions. Qualitative results are shown in Fig. 4. These results demonstrate the effectiveness in achieving robust segmentation of the proposed framework.
Tab. 2 presents the results on the synthetic benchmark DELIVER using SegFormer-B0 as the backbone. Our method achieves the highest mIoU of 46.64, representing an improvement of +6.15 over MAGIC (Zheng et al., 2024b), the previous SoTA. In individual modalities, our method achieves mIoUs of 47.11 (R), 52.17 (D), and 19.01 (L), with improvements of +14.51 (R) and +18.62 (L) compared to MAGIC. Notably, the significant gains in Event (+16.81) and LiDAR (+18.62) underscore the method’s robustness against unimodal bias. For paired modality combinations, our approach also shows strong performance. The RD, RE, and RL combinations achieve mIoUs of 60.37, 47.49, and 48.13, with notable gains of +14.47 (RE) and +15.01 (RL) over MAGIC. These results highlight the effectiveness of the model in capturing cross-modal relationships while mitigating unimodal limitations. Overall, these results validate the superiority of our method in both individual and multimodal segmentation tasks, demonstrating its ability to handle diverse data modalities and achieve robust performance.
Table 4: Ablation study on the effect of different parameters for $L_{mad}$ in our framework on MUSES dataset (Brödermann et al., 2024).
@@ 4.2 Results
Table 5: Ablation study on the effect of different parameters for $L_{umd}$ in our framework on MUSES dataset (Brödermann et al., 2024).
@@
5 Ablation Study
5.1 Effectiveness of Loss Functions
Tab. 3 shows results of different loss function combinations on the MUSES dataset (Brödermann et al., 2024). The baseline using only the supervised loss ($\mathcal{L}_{sup}$) achieves a mean mIoU of 39.29. Adding the modality-adaptive loss ($\lambda\mathcal{L}_{mad}$) improves performance, with gains in paired and combined modalities. Adding the unimodal-distillation loss ($\alpha\mathcal{L}_{umd}$) to $\mathcal{L}_{sup}$ + $\lambda\mathcal{L}_{mad}$ results in more significant gains. Notable improvements include RGB (F: +2.13), FE (+1.30), and FL (+2.58), highlighting the value of leveraging unimodal features. The full loss combination, incorporating the cross-modal distillation loss ($\beta\mathcal{L}_{cmd}$), achieves the highest mean mIoU of 40.23, with consistent improvements across all modalities. Paired modalities FL (+2.72) and FEL (+2.79) show the largest gains, while Event (E) improves slightly (+0.43). These results demonstrate the method’s robustness in handling complex multimodal interactions. Overall, each loss component plays a critical role in improving anymodal segmentation, with the full combination proving most effective in leveraging multimodal data for real-world scenarios.
5.2 Ablation on Hyper-Parameter Selection
We analyze the impact of hyper-parameters for $\mathcal{L}_{mad}$, $\mathcal{L}_{umd}$, and $\mathcal{L}_{cmd}$ in our framework. The results in Tab. 4, Tab. 5, and Tab. 6 highlight how these parameters influence performance. Tab. 4 evaluates the effect of varying $\lambda$. The baseline ($\lambda=1$) achieves a mean mIoU of 39.30. Increasing $\lambda$ improves performance, peaking at $\lambda=50$ with a mean mIoU of 40.11 (+0.81). Significant gains are observed in LiDAR (L: +2.80) and Event-LiDAR (EL: +2.09), demonstrating $\mathcal{L}_{mad}$’s effectiveness in enhancing underrepresented modalities. Further increases ($\lambda>50$) yield diminishing returns, indicating a trade-off between distillation strength and stability. Tab. 5 examines the effect of varying $\alpha$. Without $\mathcal{L}_{umd}$ (w/o), the mean mIoU is 40.11. The highest performance is achieved with $\alpha=2$, yielding a mean mIoU of 40.22 (+0.11), with notable gains in RGB (F: +2.11), FL (+2.00), and FEL (+2.29). Larger $\alpha$ values ($\alpha>5$) reduce performance, highlighting the need to balance unimodal and cross-modal contributions.
Tab. 6 studies $\mathcal{L}_{cmd}$ with varying $\beta$ while keeping $\mathcal{L}_{umd}$ fixed. The baseline ($\beta=0$) achieves a mean mIoU of 39.72. Adding $\mathcal{L}_{cmd}$ with $\beta=5$ improves the mean mIoU to 40.21 (+0.49), with strong gains in Event-LiDAR (EL: +2.04) and FL (+1.62). Larger $\beta$ values ($\beta>10$) degrade performance, suggesting overemphasis on cross-modal distillation can overshadow unimodal contributions. Optimal values of $\lambda$, $\alpha$, and $\beta$ are crucial for maximizing each loss’s benefits. While $\mathcal{L}_{mad}$ addresses unimodal biases, $\mathcal{L}_{umd}$ and $\mathcal{L}_{cmd}$ enhance both unimodal and cross-modal interactions, enabling robust anymodal segmentation.
Table 6: Ablation study on the effect of different parameters for add $L_{cmd}$ with $L_{umd}$ on MUSES dataset (Brödermann et al., 2024). w/o means the framework is only trained with $L_{mad}$ + $L_{umd}$.
@@ 5.2 Ablation on Hyper-Parameter Selection
Table 7: Discussion study on the effect of performing knowledge distillation with fused features on MUSES dataset (Brödermann et al., 2024). w/o means the framework is only trained with $L_{mad}$ + $L_{umd}$ + $L_{cmd}$.
@@
5 46.34 +0.33 16.72 -2.85 31.41 -0.72 46.79 +0.50 51.23 -0.02 33.88 -1.33 51.17 +0.03 39.65 -0.58
7 46.28 +0.27 18.01 -1.56 31.47 -0.66 46.76 +0.47 51.05 -0.20 33.99 -1.22 50.97 -0.17 39.79 -0.44
10 46.10 +0.09 15.61 -3.96 31.39 -0.74 46.55 +0.26 51.23 -0.02 33.59 -1.62 51.14 +0.00 39.37 -0.86
5.3 Rationality of Unimodal Distillation
The effectiveness of $\mathcal{L}_{umd}$ is demonstrated in Tab. 5. Adding $\mathcal{L}_{umd}$ with $\alpha=1$ improves single-modality performance, particularly for RGB ($F:+0.83$) and paired modalities such as FL ($+0.42$). The best performance for single-modality tasks is observed at $\alpha=10$, where the mean mIoU for RGB increases to 46.17 ($+2.46$), and FL achieves 51.17 ($+2.19$). Obviously, $\mathcal{L}_{umd}$ effectively improves single-modality segmentation performance by facilitating knowledge transfer within each modality. However, the results also highlight a trade-off: while unimodal distillation enhances performance for individual modalities, it may compromise the model’s ability to handle complex multimodal combinations. Careful tuning of $\alpha$ is therefore critical to achieving a balance between single-modality and multimodal segmentation performance.
@@ 5.3 Rationality of Unimodal Distillation
5.4 Why not KD between Fused Features?
To explore the possibility of transferring knowledge between fused features from teacher to student models, we conduct experiments on the MUSES dataset, leveraging the fusion method described in (Zheng et al., 2024b). The results are summarized in Tab. 7. It is evident that applying knowledge distillation (KD) directly on fused features causes a notable performance drop across most metrics. For instance, compared to the baseline without KD on fused features (w/o), which achieves a mean mIoU of 40.23, all settings with KD exhibit reduced performance. At $\lambda=1$, the mean mIoU drops to 39.77 (-0.46), while higher $\lambda$ values exacerbate the decline, with $\lambda=10$ resulting in a mean mIoU of 39.37 (-0.86). These results highlight the inefficacy of such distillation methods. Examining individual modalities further reveals this trend. For the Event (E) and LiDAR (L) modalities, performance consistently degrades as $\lambda$ increases, with Event mIoU decreasing by up to -3.96 at $\lambda=10$. Paired and fused modalities also show minimal improvement or slight degradation, such as FEL achieving a marginal gain of +0.03 at $\lambda=5$ but regressing at $\lambda=10$. The results suggest that directly applying knowledge distillation between fused features fails to effectively guide the student model. Instead of transferring meaningful knowledge, the fused features introduce noise and misalignment, limiting their utility in this context. This reinforces the importance of designing specialized mechanisms for knowledge distillation, targeting individual or structured features rather than indiscriminately fused representations.

### v2

@@ F E L FE FL EL FEL
CMX [5] T-ITS 2023 2.52 2.35 3.01 41.15 41.25 2.56 42.27 19.30
CMNeXt [17] CVPR 2023 FEL 3.50 2.77 2.64 6.63 10.28 3.14 46.66 10.80
MAGIC [7] ECCV 2024 43.22 2.68 22.95 43.51 49.05 22.98 49.02 33.34
Any2Seg [6] ECCV 2024 44.40 3.17 22.33 44.51 49.96 22.63 50.00 33.86
@@ 3.3 Modality-agnostic Distillation
The total loss for training the anymodal student segmentor combines the supervised objective with multiple distillation terms, formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{sup}+\lambda_{\text{mad}}\mathcal{L}_{mad}+\alpha\mathcal{L}_{umd}+\beta\mathcal{L}_{cmd},$ (8)
where $\lambda_{\text{mad}}$, $\alpha$, and $\beta$ are weighting coefficients for the modality-adaptive, unimodal, and cross-modal distillation losses, respectively. The supervised loss $\mathcal{L}_{sup}$ serves as the primary training signal and is left unweighted for clarity, with all distillation losses scaled relative to it.
Table 2: Results of anymodal semantic segmentation validation with four modalities (R: RGB, D: Depth, E: Event, L: LiDAR) on DELIVER using SegFormer-B0 as backbone.
@@
CMNeXt [17] 0.86 0.49 0.66 0.37 47.06 9.97 13.75 2.63 1.73 2.85 59.03 59.18 14.73 59.18 39.07 20.77
MAGIC [7] 32.60 55.06 0.52 0.39 63.32 33.02 33.12 55.16 55.17 0.26 63.37 63.36 33.32 55.26 63.40 40.49
@@
Table 3: Ablation study of different loss combinations on MUSES dataset [18].
@@
Table 4: Ablation study on the effect of different parameters for $L_{mad}$ in our framework on MUSES dataset [18]. More results can be found in Table 11.
@@
4 Experiments
Experimental Setup. We evaluate our method on synthetic and real-world multi-sensor datasets. The MUSES dataset [18], recorded in Switzerland, includes driving sequences designed to address challenges from adverse visual conditions. It features multi-sensor data from a high-resolution frame camera, an event camera, and MEMS LiDAR, enhancing annotation quality and supporting robust multimodal semantic segmentation. Each sequence is annotated with 2D panoptic labels for accurate ground truth. The DELIVER dataset [17] includes RGB, depth, LiDAR, and event data across 25 semantic categories, covering various environmental conditions and sensor failures for thorough evaluations. We follow the official data processing and split protocols. More implementation details can be found in Sec. A.1
4.1 Results
As shown in Tab. 1, our method achieves the highest mIoU of 40.23, surpassing all state-of-the-art (SoTA) baselines. CMX [5] and CMNeXt [17] perform poorly (mIoU: 19.30% and 10.80%), due to over-reliance on RGB. Similarly, MAGIC [7] and Any2Seg [6] depend heavily on depth, leading to large performance drops when it’s absent—exposing their unimodal bias. In contrast, our method shows strong, balanced performance across all modalities: RGB (46.01%), Event (19.57%), and LiDAR (32.13%), as well as paired combinations: FE (46.29%), FL (51.25%), EL (35.21%), and FEL (51.14%), demonstrating robust cross-modal learning. Significant gains in Event (+16.40%) and FL (+12.58%) further highlight the ability to handle challenging or sparse modalities. Fig. 3 provides qualitative comparisons. Tab. 2 reports results on the synthetic DELIVER benchmark using SegFormer-B0. Our method achieves a top mIoU of 46.64%, outperforming MAGIC [7] by +6.15%. On individual modalities, it achieves 47.11% (R), 52.17% (D), and 19.01% (L), with +14.51% (R) and +18.62% (L) gains over MAGIC. Event and LiDAR improvements further confirm robustness to unimodal bias. For paired modalities, RD, RE, and RL achieve mIoUs of 60.37%, 47.49%, and 48.13%, with notable gains over MAGIC (+14.47% RE, +15.01% RL). These results demonstrate our model’s strength in capturing cross-modal dependencies while remaining resilient to missing inputs. Overall, the results confirm the effectiveness of our method in both unimodal and multimodal settings, offering strong generalization across diverse sensor combinations.
Table 5: Ablation on different parameters for $L_{umd}$ in our framework on MUSES dataset [18].
@@ 4.1 Results
Table 6: Ablation on different parameters for add $L_{cmd}$ with $L_{umd}$ on MUSES dataset [18]. w/o means the framework is only trained with $L_{mad}$ + $L_{umd}$. More results can be found in Table. 12.
@@
5 Ablation Study
Effectiveness of Loss Functions Tab. 3 reports results of different loss combinations on the MUSES dataset [18]. Using only supervised loss ($\mathcal{L}_{sup}$) yields a mean mIoU of 39.29%. Adding $\lambda\mathcal{L}_{mad}$ improves performance across paired and combined modalities. Further including $\alpha\mathcal{L}_{umd}$ brings notable gains, especially in F (+2.13%), FE (+1.30%), and FL (+2.58%), confirming the benefit of unimodal knowledge distillation. The full combination with $\beta\mathcal{L}_{cmd}$ achieves the best performance (mean mIoU 40.23%), with broad improvements, particularly in FL (+2.72%) and FEL (+2.79%). Even E improves slightly (+0.43%), showing robustness across modalities. In sum, each loss contributes to better segmentation, and the full loss design is most effective for capturing complex multimodal interactions.
Ablation on Teacher Model and Fusion Strategy We compare our PML with fusion-based teacher models MAGIC and CMNeXt to assess their impact on student performance. Unlike CMNeXt’s fixed fusion architecture, PML enables both unimodal and cross-modal distillation via parallel modality learning, providing more diverse and semantically aligned supervision.
Table 7: Ablation on teacher model selection. Results reported as mIoU (%).
Teacher F E L FE FL EL FEL Mean
MAGIC 43.87 13.22 30.90 43.91 48.41 33.68 47.75 37.39 (-2.84)
CMNeXt 43.80 10.79 23.15 48.34 43.97 33.52 47.64 35.89 (-4.34)
PML (Ours) 46.01 19.57 32.13 46.29 51.25 35.21 51.14 40.23
Results in Tab. 7 show that distilling from fused features, as in MAGIC and CMNeXt, degrades performance. CMNeXt yields the largest drop (–4.34% mIoU), indicating that rigid fusion limits generalization. In contrast, PML achieves the best performance by preserving both modality-specific and shared representations. These results confirm PML’s flexibility and effectiveness as a general-purpose teacher for anymodal segmentation.
Ablation on Hyper-Parameter Selection We study the effect of hyper-parameters $\lambda$, $\alpha$, and $\beta$ for $\mathcal{L}_{mad}$, $\mathcal{L}_{umd}$, and $\mathcal{L}_{cmd}$, respectively (Tab. 46). Increasing $\lambda$ improves performance up to a point, particularly benefiting underrepresented modalities, before plateauing due to diminishing returns. Varying $\alpha$ reveals that moderate values enhance unimodal contributions, while overly large values lead to degradation. Adding $\mathcal{L}_{cmd}$ with an appropriate $\beta$ further improves results, but excessive weighting can suppress unimodal learning.
Table 8: Discussion study on the effect of performing KD with fused features on MUSES dataset [18]. w/o means the framework is only trained with $L_{mad}$ + $L_{umd}$ + $L_{cmd}$.
@@ 5 Ablation Study
Rationality of Unimodal Distillation The effectiveness of $\mathcal{L}_{umd}$ is demonstrated in Tab. 5. Adding $\mathcal{L}_{umd}$ with $\alpha=1$ improves single-modality performance, particularly for RGB ($F:+0.83\%$) and paired modalities such as FL ($+0.42\%$). The best performance for single-modality tasks is observed at $\alpha=10$, where the mean mIoU for RGB increases to 46.17 ($+2.46\%$), and FL achieves 51.17% ($+2.19\%$). Obviously, $\mathcal{L}_{umd}$ effectively improves single-modality segmentation performance by facilitating knowledge transfer within each modality. However, the results also highlight a trade-off: while unimodal distillation enhances performance for individual modalities, it may compromise the model’s ability to handle complex multimodal combinations. Careful tuning of $\alpha$ is therefore critical to achieving a balance between single-modality and multimodal segmentation performance.
@@
Why not doing KD between Fused Features? To explore the possibility of transferring knowledge between fused features from teacher to student models, we conduct experiments on the MUSES dataset, leveraging the fusion method described in [7]. The results are summarized in Tab. 8. It is evident that applying knowledge distillation (KD) directly on fused features causes a notable performance drop across most metrics. For instance, compared to the baseline without KD on fused features (w/o), which achieves a mean mIoU of 40.23%, all settings with KD exhibit reduced performance. At $\lambda=1$, the mean mIoU drops to 39.77% (-0.46%), while higher $\lambda$ values exacerbate the decline, with $\lambda=10$ resulting in a mean mIoU of 39.37% (-0.86%). These results highlight the inefficacy of such distillation methods. Examining individual modalities further reveals this trend. For the Event (E) and LiDAR (L) modalities, performance consistently degrades as $\lambda$ increases, with Event mIoU decreasing by up to -3.96% at $\lambda=10$. Paired and fused modalities also show minimal improvement or slight degradation, such as FEL achieving a marginal gain of +0.03% at $\lambda=5$ but regressing at $\lambda=10$. The results suggest that directly applying knowledge distillation between fused features fails to effectively guide the student model. Instead of transferring meaningful knowledge, the fused features introduce noise and misalignment, limiting their utility in this context. This reinforces the importance of designing specialized mechanisms for knowledge distillation, targeting individual or structured features rather than indiscriminately fused representations.
t-SNE Visualization Fig. 4 presents t-SNE visualizations of multimodal features, including RGB, Depth, Event, LiDAR, and the features under our AnySeg framework. Individual modality plots reveal distinct clusters, reflecting semantic separability. RGB and Depth exhibit relatively compact clusters, indicating strong discriminative power, whereas Event and LiDAR show more dispersed and overlapping clusters, highlighting weaker performance when used independently. The learned feature spaces, particularly AnySeg-RDL and AnySeg-RDEL, show notable improvements in cluster compactness and separability. AnySeg-RDEL, integrating RGB, Depth, Event, and LiDAR, achieves the most coherent and well-separated clusters, demonstrating the framework’s robustness in leveraging complementary information across modalities. These results underscore AnySeg’s effectiveness in addressing the limitations of individual modalities, achieving robust multi-modal feature representation, and enhancing segmentation performance through cross-modal fusion.
Table 9: Performance under different conditions in the DELIVER with RGB-D modalities.
Metric All Cloud Fog Night Rain Sun M.B. O.E. U.E. L.J. E.L.
mIoU (%) 55.99 57.94 55.02 54.64 56.24 56.82 54.73 53.71 53.15 54.66 54.18
mF1 (%) 66.97 68.80 65.33 66.16 67.43 67.46 64.88 64.59 64.10 64.51 65.04
mAcc (%) 64.58 65.83 64.51 63.62 64.91 64.41 63.82 61.88 62.06 62.46 62.65
Cross-Sensor Generalization under Scene Degradations. We test AnySeg’s performance on the DELIVER using RGB-D under various challenging conditions. As shown in Tab. 9, AnySeg maintains consistent performance across diverse sensor and environmental degradations. The model handles cloud, rain, and sun conditions well, with only minor drops under more severe challenges like motion blur, extreme exposure, and sensor noise (e.g., lidar jitter, event resolution). Despite these disruptions, the model achieves over 53% mIoU in all cases, demonstrating strong generalization.
Figure 4: TSNE visualization of multi-modal features (RGB-R, Depth-D, Event-E, and LiDAR-L) and the learned features of our AnySeg framework.